OpenStudy (anonymous):

Write an equation in slope intercept form for the line that passes through (-3,8) and (4,-2)

OpenStudy (abb0t):

\[slope = m = \frac{ y_2-y_1 }{ x_2-x_1 }\] then\[(y-y_0)=m(x-x_0)\]

OpenStudy (anonymous):

good enough Thanx abbot